OpenStudy (anonymous):

Consider the following reaction: 2 H2O(g) + 2 SO2(g) ⇌ 2 H2S(g) + 3 O2(g) A reaction mixture initially contains 2.8 M H2O and 2.6 M SO2. Determine the equilibrium concentration of H2S if Kc for the reaction at this temperature is 1.3 ×10^‐6

OpenStudy (abb0t):

start by setting up your equilibrium equation: \(\sf \color{red}{K_c} = \frac{products}{reactants}\)

OpenStudy (abb0t):

In your case, you would use: \(\sf \color{red}{K_c} = \frac{[H_2S]^2[O_2]^3}{[SO_2]^2[H_2O]^2}\)

OpenStudy (abb0t):

1.3 \(\times\) 10\(^{-6}\) = \(\sf \color{red}{K_c} = \frac{[H_2S]^2[O_2]^3}{[SO_2]^2[H_2O]^2}\)